OpenStudy (anonymous):

Emily went to a toy store that was selling one Nintendo DS game and one Wii game for a total of $70. Emily bought 3 DS games and 2 Wii games and spent $160. What was the price of one Wii game? $20 $30 $40 $50

OpenStudy (anonymous):

$20?

OpenStudy (anonymous):

Let x denote price of Nintendo Ds game and y denote Wii game x + y = 70 ---- equation 1 3x + 2y = 160 ---- equation 2 y = 70 - x ---- equation 3 Plugging y from equation 3 into equation 2 3x + 2 (70 -x) = 160 3x + 140 - 2x = 160 x = 160 - 140 x = 20 Price of Nintendo Ds game = 20$ and now we find wii game price Using x in equation 1 x + y = 70 Plugging x = 20 we get 20 + y = 70 y = 70 - 20 y = 50 So the price of Nintendo Wii game is 50$